OpenStudy (anonymous):

how do i factor xsquard-x-2=0

OpenStudy (gw2011):

x^2-x-2=0 (x-2)(x+1)

OpenStudy (anonymous):

how

OpenStudy (gw2011):

If you look at x^2 which is the first term, the factors are (x) times (x) Now look at the last term of -2. The factors are either -2 and +1 or +2 and -1. You have to look at the middle term which is -x to determine which factors will work and then you will see that the factors of -2 and +1 will give you -2x+x which equals -x. Recapping: 1) Find the factors of the first term 2) Find the factors of the last term 3) The combination of the factors of the first term and the last term must equal the middle term. For example, if you look at the answer to the above problem, you would have to multiply the middle terms (which in this case is( -2)(x)=-2x) and add it to the outside terms (which in this case is (x)(1)=x) and the result will be -2x+x=-x which is the middle term in the problem.